] a. The equation of Lineweaver-Burk double-reciprocal plot of caffeine dehydrogenase-catalyzed reaction is   y = 12x...

  1. ] a. The equation of Lineweaver-Burk double-reciprocal plot of caffeine dehydrogenase-catalyzed reaction is   y = 12x + 3. Calculate Vmax (mmol/s) and Km (mmol/L).

b. [5 points] Estimate V for caffeine concentration of 400 mmol/L.

c. [10 points] The enzyme caffeine dehydrogenase (Cdh) is an inducible quinone-dependent oxidoreductase. Describe how the addition of caffeine into the culture medium will be detected and transcriptionally regulated by the two-component system in Pseudomonas sp. CBB1.
Answer #1

(a)

We have, from the Lineweaver-Burke plot:

1 Km 1 1 [V] – Vmar [51 + Vmar

Now, when Km/Vmax[S] = 0, 1/[V] = 1/Vmax

We have,

y = 12x + 3

When x = 0;
y = 3
=> 1/[V] = 3
=> 1/Vmax = 3
=> Vmax = 1/3 = 0.33 mmol/s

When y = 0, that is 1/[V] = 0;

12x = -3
=> x = -3/12 = -0.25

=> -1/Km = -0.25
=> 1/Km = 0.25
=> Km = 4 mmol/L

Therefore,
Vmax = 0.33 mmol/s ; Km = 4 mmol/L

(b) From the Lineweaver-Burke equation:

1 km 1. 1 v = Vmar [S] * Vmar
14 - 4 - 1 - 1,1 - 1 100 101 0.33 x 400*0.33 330.33333333

== 3.0606

=> V = 0.3267

The V of the reaction at [S] = 400 mmol/L is 0.3267 mmol/s